# Bridge indices of spatial graphs and diagram colorings

### v1

@@ 3 Main result
[…]
A sequence of coloring moves can be notated as $(A_{0},f_{0})\rightarrow(A_{1},f_{1})\rightarrow\cdots\rightarrow(A_{m},f_{m})$, where $f_{i}$ is a function from the subset of strands $A_{i}$ to a finite set of $k$ elements, which are the colors. To elaborate, $(A_{0},f_{0})$ is the stage where only the seeds have been colored with the $k$ colors. The strands that used to be attached to a pod will get repeated colors. The final stage $(A_{m},f_{m})$ is when all the strands get colored.
In the proof of our main result, we embed each strand at a certain height dictated by the coloring moves. Thus, there is a function $h\colon\{s_{1},s_{2},\cdots,s_{n}\}\rightarrow\mathbb{R}$ from the set of strands that records the height of each strand $s_{i}$.
Lemma 3.1.
At any coloring stage, each color in the diagram is either (1) a connected arc if the seed corresponding to that color is an arc or (2) $n$ disjoint arcs if the seed corresponding to that color is a pod of degree $n$.
Proof.
We verify this claim by induction on $u$, where $u$ is the coloring stage. For the base case $u=0,$ this is the stage where no coloring move has been applied, so only the seeds are colored. The claim immediately holds by the way the seeds are defined. Now assume that the claim holds for $u<t$ and consider the stage $u=t>0$. By the definition of coloring move, there is a unique strand $s_{i}$ adjacent to a strand in $A_{t-1}$ such that $A_{t}\backslash A_{t-1}=\{s_{i}\}$. Thus, $f_{t}^{-1}(c)=f_{t-1}^{-1}(c)\cup\{s_{i}\}$ for a particular color $c.$ By the inductive hypothesis, we have that $f_{t-1}^{-1}(c)$ is either $n$ arcs if the seed corresponding to that color is a pod of degree $n$, or a connected arc if the seed corresponding to that color is an arc. Adding one arc $s_{i}$ adjacent to some arcs in $f_{t-1}^{-1}(c)$ preserves the property stated in the hypothesis. ∎
Lemma 3.2.
Suppose that a nontrivial spatial graph diagram $D$ is $k$-colorable via the sequence $(A_{0},f_{0})\rightarrow(A_{1},f_{1})\rightarrow\cdots\rightarrow(A_{n},f_{n})$. At a crossing $c$, let $s_{p},s_{q},$ and $s_{r}$ denote an incoming understrand, an outgoing understrand, and an overstrand, respectively. If the understrands receive the same color $j$ at stage $n$, then one of the following situations occurs:
  1. (1)
    the overstrand is not colored last,
  2. (2)
    the strands with the same color correspond to an unknotted component and $c$ is the unique crossing with the property that $h(s_{r})\leq\min\{h(s_{q}),h(s_{p})\}$, or
  3. (3)
    the strands with the same color correspond to self-loops (that is, edges which connect a vertex to itself) in the graph connected to a pod and $c$ is the unique crossing in a self-loop with the property that $h(s_{r})\leq\min\{h(s_{q}),h(s_{p})\}$.
Proof.
Suppose that $h(s_{r})\leq\min\{h(s_{q}),h(s_{p})\}$. First, suppose that $s_{p}=s_{q}$. Then, the lift of $s_{p}=s_{q}$ is an unknotted component in 3-space. Furthermore, the only way to obtain a Wirtinger colorable diagram is to make $s_{p}=s_{q}$ a seed strand. In this case, we get the conclusion (2). The uniqueness of the crossing with this property is automatic since this unknotted component only has one crossing with the rest of the diagram in which it is the understrand.
Now we may suppose that $s_{p}\neq s_{q}$. Without loss of generality, suppose that $h(s_{q})\leq h(s_{p})$. By Lemma 3.1, two cases can happen.
Case 1 (the color $j$ forms a connected arc because the corresponding seed is an arc): The strand $s_{q}$ receives its color from an adjacent strand $s_{q_{1}}$ by a coloring move. There is also a stage in which $s_{q_{1}}$ receives its color from an adjacent strand $s_{q_{2}}$ by a coloring move. We can continue in this fashion to see that there is a stage in which $s_{q_{k}}$ receives its color from a seed. This seed may or may not be $s_{p}$. If it is $s_{p}$, then we have found an unknotted component. If not, then we can do the same backtracking process starting at $s_{p}$ to reach the seed for the color $j,$ forming an unknotted component. Note that the existence of another crossing with the property that $h(s_{r})\leq\min\{h(s_{q}),h(s_{p})\}$ implies that there were at least two seed arcs. However, repeated colors are only allowed for pod seeds.
Case 2 (the color $j$ consists of multiple arcs converging to a pod): This case is depicted in Figure 4. Essentially, we use a backtracking argument starting at $s_{q}$ as in Case 1 until we reach a strand in the truncated diagram with an open end. Backtracking from $s_{p}$ will also end at the same pod. This creates the desired loop in the graph. The existence of another crossing with the property that $h(s_{r})\leq\min\{h(s_{q}),h(s_{p})\}$ is only possible if multiple pod seeds have the same color. However, our coloring rule dictates that there is only one arc or one pod such that all strands connected to it receive the color $j$. ∎
Figure 4: The crossing involving the black (uncolored) strand has the property that the the overstrand $s_{r}$ is not colored yet at this stage, but the understrands are colored with the same color. Here the vertex is put back into the truncated diagram for reference.
Remark 3.3.
The uniqueness of a crossing for a color with the property that $h(s_{r})\leq\min\{h(s_{q}),h(s_{p})\}$ will be used to ensure that there are the correct number of places to put a local minimum in the proof of Theorem 1.1.
@@
Given a spatial graph, suppose that its bridge index $r$ is realized by a diagram $D$, and without loss of generality, let the number of upper pods and maxima be less than or equal to the number of lower pods and minima. Think of $D$ as lying in the $xy$-plane. We can pull the upper pods and maxima to the same height $y=a$, and the lower pods and minima to the same height $y=b$. After making the $r$ upper pods and maxima in the diagram seeds, the colorings can be extended to the entire diagram using only the coloring moves. To see this, observe that after removing the maxima, minima, and vertices of degree at least three at the pods, we obtain a braid diagram between $y=a$ and $y=b$. Once the seeds are chosen, all of the strands connected to the top endpoints of the braid diagram are colored. If the braid is slightly perturbed so that each crossing is at a distinct height, it is easy to see that the coloring can be extended at every crossing (see Figure 5 and its caption). We can also require that the lower pods are colored last. Therefore, the Wirtinger number of $D$ is $r$. ∎
Figure 5: In a block of a braid diagram that contains only one crossing, there is only one strand that is not connected to the top dashed horizontal boundary. Such a strand is colored black to signify that it has not been colored yet. Thus, if the strands connected to the top are colored, then we just need one coloring move for each block of braid diagram.
@@
By Proposition 3.4 we have that the Wirtinger number is at most the bridge index. For the other inequality, suppose we have a spatial graph with Wirtinger number $n$ which is realized by a diagram $D$. We will construct an embedding of our graph in $\mathbb{R}^{3}$ with bridge index $n$. Refer to Figure 6 throughout the proof for an example of how the embedding works.
We embed the strand that receives color at stage $j$ in the plane $z=-j$ in $\mathbb{R}^{3}$. Here, strands belonging to the same seed pod are on the same level. If the graph has $m$ strands, we embed the diagram $D$ itself in the level $z=-m-5$. The exact level does not matter; the spatial graph will sit at or above level $z=-m$ by construction, and we just wish to place the diagram below so that the graph projects down to the diagram. We denote the lift of the strand $s_{i}$ in the diagram to the spatial graph by $\widetilde{s_{i}}$.
At the moment, the embedded strands are all disconnected. Note that the disconnected places correspond to either a crossing in the diagram or the free ends coming from removing the vertices of the pods. To form an embedding of our spatial graph, we will connect up the strands crossing-by-crossing in such a way that our embedding has $k$ upper pods and $n-k$ local maxima, which will project down to seeds in the diagram. Then we will add in the vertices of the pods. For each crossing, this connection will happen in a cylinder $B^{2}_{\varepsilon}(c)\times\mathbb{R}$, where $B^{2}_{\varepsilon}(c)\subset\{z=-m-5\}$ is a open 2-disk neighborhood of the crossing $c$ with radius $\varepsilon$ chosen small enough so that it contains only the overstrand $s_{k}$ and the two adjacent understrands $s_{i}$ and $s_{j}.$ Each crossing falls into one of the following two cases.
Case 1: Assume that at the end of the coloring process, the color of $s_{i}$ is not the same as the color of $s_{j}$. In this case, we connect one end of the lift of $s_{i}$ to an end of the lift of $s_{j}$ with an arc $s_{ij}$ so that exactly one local minimum is created. We cannot use a monotonic arc as the height of the overstrand is too low. More precisely, $s_{ij}$ is constructed as a union of two arcs. One of the arcs starts from $\widetilde{s_{i}}$, and goes monotonically down to a point $x_{ij}$, where $x_{ij}$ projects down to $c$ in the diagram. The other arc goes from $x_{ij}$ monotonically up to $\widetilde{s_{j}}$. We note that there are many choices for where to put $x_{ij}$. For instance, putting $x_{ij}$ at the height $z=-m-4$ will work. In Figure 6, the point $x_{ij}$ is the local minimum of the green arc.
Figure 6: A schematic showing that the unique multi-colored crossing (highlighted in gray) corresponds to connecting up the disjoint strands with a new strand (colored green) with a single local minimum.
Case 2: Assume that at the end of the coloring process, the color of $s_{i}$ is the same as the color of $s_{j}$. In this case, by Lemma 3.2, there are two subcases.
Subcase 2.1: Suppose it holds that $h(s_{k})>\min\{h(s_{i}),h(s_{j})\}$. Without loss of generality, suppose that $\widetilde{s_{i}}$ is lower than $\widetilde{s_{j}}$. For this case, the connection arc can be created without introducing a local maximum or a local minimum because the overstrand is not colored last. Formally, we construct $s_{ij}$ as a monotonically decreasing arc from an endpoint of $\widetilde{s_{j}}$ to an endpoint of $\widetilde{s_{i}}$ contained in $B^{2}_{\varepsilon}(c)\times\mathbb{R}$ so that a point on $s_{ij}$ orthogonally projects to the crossing $c$. For a concrete demonstration, please consult Figure 6, where every crossing except the one highlighted with the grey oval falls into this subcase.
Subcase 2.2: Assume that at the end of the coloring process, the color of $s_{i}$ is the same as the color of $s_{j}$ and $h(s_{k})\leq\min\{h(s_{i}),h(s_{j})\}$. Then by Lemma 3.2, we either get an unknotted component or a self-loop (that is, an edge which connects a vertex to itself) in our spatial graph.
In the former case, Lemma 3.2 implies that there is exactly only one crossing $c$ with this property such that the overstrand receives the color last. We then connect the strand $\widetilde{s_{i}}$ to $\widetilde{s_{j}}$ by two arcs as in Case 1 so that this crossing $c$ corresponds to exactly one local minimum in the lift. In the latter case, Lemma 3.2 implies that there is exactly one crossing $c_{i}$ for each self-loop $i$. We also connect the strand $\widetilde{s_{i}}$ to $\widetilde{s_{j}}$ by two arcs as in Case 1 so that each crossing $c_{i}$ corresponds to exactly one local minimum in the lift.
After all open ends are joined up at each crossing, we can add vertices to the levels containing multiple strands. After a slight perturbation, these vertices correspond to upper pods. The lifts of the other seed strands, where each seed was one connected arc in the initial coloring stage, can be perturbed slightly to be local maxima. The multi-color crossings lift to local minima.
The remaining disconnected places in the lift project down to vertices of non-seed pods in the diagram. For such a vertex $v$, let the lifts of the strands be $\{\widetilde{s_{i_{1}}},\widetilde{s_{i_{2}}},\ldots,\widetilde{s_{i_{d}}}\}$, where $d$ is the degree of the corresponding non-seed pod. To create our desired embedding, we place a point $\widetilde{v}$ at the free end of the lowest arc, say $s_{i_{d}}$ without loss of generality. Then, add an arc from the free end of $\{\widetilde{s_{i_{1}}},\widetilde{s_{i_{2}}},\ldots,\widetilde{s_{i_{d-2}}},\widetilde{s_{i_{d-1}}}\}$ monotonically down to $\widetilde{v}$. Each such $\widetilde{v}$ will correspond to lower pods after a slight perturbation. This completes our embedding: after possibly some additional slight perturbations, we have put our spatial graph in bridge position with bridge index $n$. ∎
@@ 4 Lower bounds
The following techniques are computable quantities that bound the bridge index of graphs from below. In the setting of links, we know by experimentation that depending on the link types, one method may be more effective than the others [vo2024learning]. That is, in some instances the computer code finds enough quandle colorings, but not a rank bound, and vice versa.
@@ 4.1 Quandle colorings
In [fleming2007virtual, niebrzydowski2010coloring] the authors generalized knot quandles to spatial graphs.
@@
A consistent labeling [^1] Various authors also use the term quandle coloring, $X$-coloring, or simply coloring for this of a diagram by a quandle is an assignment of a quandle element to each strand, such that strands connected to the same vertex receive the same label, and the relation in Figure 7 is satisfied at each crossing. It can be shown that for knots and links, the number of consistent labelings is unchanged by the Reidemeister moves. For general spatial graphs, however, this is no longer true.
@@
Figure 8: For a quandle coloring to work, the strands near a crossing need to all receive the same color or all distinct colors. On the left, the strands are disjoint and the strands adjacent to a pod receive one color. If the other strand receives a different color, after we perform a move that does not change the graph type, we are forced to have a crossing that uses only two colors.
@@
This motivates the following definition.
@@
We remark that in [niebrzydowski2010coloring], the authors allow edges coming out of the same vertex to have different labels, but for our lower bound of the bridge index $\beta$, we will just consider the case where the elements at the edges around the same vertex are the same.
@@
Let $X$ be a finite $n$-quandle of order $|X|.$ Let $Col_{X}(G)$ denote the coloring number of the spatial graph $G$ by $X$. Then $Col_{X}(G)\leq|X|^{\beta(G)}$. In other words,
$\log_{|X|}Col_{X}(G)\leq\beta(G).$
@@
Some elements of this proof appeared in Proposition 3.4, but instead of just using colors, we are coloring with quandle elements. There are $|X|$ possibilities to assign an element of $X$ to an upper pod or a local maximum. In a bridge index minimizing diagram, there are $|X|^{\beta(G)}$ ways to assign elements of $X$ to the upper pods and local maxima. This assignment determines all of the other labels in the diagram as argued in Proposition 3.4. ∎
We now provide a lower bound for the bridge index of vertex sums of spatial graphs (recall Section 2.1) in terms of quandles, which will be useful to us in Section 5. The argument is based on [clark2016quandle]. The key is the existence of quandles called homogeneous quandles, which have the property that for any quandle elements $x,y,$ there is an automorphism $h$ such that $h(x)=y.$ Note that the map $f_{y}$ sending $x$ to $x\triangleright y$ in the second axiom of Definition 4.1 is an example of an automorphism by the way the axiom is stated. The reader can check from the table before Definition 4.2 that the Alexander quandle of order four is homogeneous.
@@
Let $X$ be a homegeneous $n$-quandle, and $G\#_{n}G^{\prime}$ be the $n$-valent vertex sum of spatial graphs $G$ and $G^{\prime}$. Then
@@
$\log_{|X|}Col_{X}(G)+\log_{|X|}Col_{X}(G^{\prime})-1\leq\beta(G\#_{n}G^{\prime}).$
@@
The number of colorings where a fixed color $x$ is assigned to the strands adjacent to a vertex $v$ is precisely $Col_{X}(G)/|X|$. To see this, observe that the number of colorings $Col_{X}(G)=\sum_{y}Col_{X}(G,y)$, where $Col_{X}(G,y)$ denotes the number of colorings such that the strands near $v$ get the label $y$. We can get another coloring in which the strands adjacent to $v$ are colored $z\neq y$ by applying an automorphism $h$ such that $h(y)=z,$ which exists by the homogeneity assumption. Thus, $Col_{X}(G)=|X|Col_{X}(G,y)$.

### v2

@@ 3 Main result
This section is dedicated to the proof of Theorem 1.1. We start with two relevant definitions.
Definition 3.1.
A spatial graph diagram $D$ realizes the bridge index for the spatial graph $G$ it represents if a circle can be drawn in the plane in which the diagram sits that splits the diagram into diagrams for two trivial tangles, the tangle diagrams intersect the circle in $b$ points, and the bridge index of $G$ is $\tfrac{1}{2}b$.
Similarly, a spatial graph diagram $D$ realizes the Wirtinger number for the spatial graph $G$ it represents if there exist seeds which completely color $D$ and whose weights sum to $2k$, and the Wirtinger number of $G$ is $k$.
One direction of our main result is more immediate; we consider this direction in the following proposition. We phrase our proofs in this section in terms of spatial graphs in $\mathbb{R}^{3}$ instead of $S^{3}=\mathbb{R}^{3}\cup\{\infty\},$ but this does not make the result less general since the isotopy classes of spatial graphs in $S^{3}$ and $\mathbb{R}^{3}$ are the same: the point at infinity $\{\infty\}$ can be taken to be disjoint from a spatial graph and its isotopies.
@@
Let $D$ be a diagram of $G$ that realizes the bridge index. There exists a circle $C$ that intersects $D$ in $\beta(G)$ points that separates $D$ into diagrams of trivial tangles $\tau_{1}$ and $\tau_{2}$. We embed $G$ in $\mathbb{R}^{3}$ such that it projects to $D$ on the $xy$-plane, and adjust the height ($z$-coordinate) of pieces of the graph as follows. After a slight deformation, we can assume that the crossings and vertices lie in different levels. Furthermore, we can arrange the local maxima and vertices to be higher than crossings, and arrange the local minima and vertices to be lower than crossings. We then choose as seeds all of the strands that intersect $C$. The colorings at these seeds extend to the entire diagram because the colorings can be extended one crossing at a time as shown in Figure 7. Afterwards, all the free ends of every tree will be colored. ∎
Figure 7: Extending the coloring from all the strands that intersect a projected bridge sphere to all the strands in the trivial tangle $\tau_{2}$. Extending to $\tau_{1}$ is similar.
Now we prove our main result. We define a “doubling map” $P\colon\mathscr{G}\rightarrow\mathscr{L}$ from the set $\mathscr{G}$ of diagrams (in $\mathbb{R}^{2}$) of spatial graphs in $S^{3}$ to the set $\mathscr{L}$ of diagrams (in $\mathbb{R}^{2}$) of links in $S^{3}$, where $P$ doubles each edge, removes a neighborhood of each vertex, and connects up adjacent arcs as shown in Figure 8. We will use our doubling map to reduce the case of spatial graphs to the case of links, and then apply the results of [Bla+20a].
Figure 8: The effect of the doubling map $P$ on a spatial graph diagram.
@@
By Proposition 3.2 we have that the Wirtinger number is at most the bridge index.
For the other inequality, suppose we have a spatial graph $G$ with Wirtinger number $k$ which is realized by a diagram $D_{G}$. Apply $P$ to get $P(D_{G})\vcentcolon=D_{L}$, which is the diagram of a link $L$. Pick seeds $\{s_{1},\dots,s_{n}\}$ in $D_{G}$ that are sufficient to color the entire diagram in such a way which yields the Wirtinger number $k$. These seeds then induce seeds in $D_{L}$, but with the rule that different strands in $D_{L}$ coming from the same seed in $D_{G}$ must count as different seeds and thus be colored differently, in accordance with the usual procedure for determining seeds for links; see Figure 9. The number of seeds in $D_{L}$ coming from the $n$ seeds $\{s_{1},\dots,s_{n}\}$ in $D_{G}$ will be $\sum_{i=1}^{m}w_{i}=2k$, where $w_{i}$ is the weight of the seed $s_{i}$ in $D_{G}$.
Figure 9: The three seeds in $D_{G}$ on the left induce seven seeds in $D_{L}$ on the right.
We then claim that these induced seeds in $D_{L}$ are sufficient to color the whole diagram. This follows from an analysis of the effect of the map $P$ on the coloring move plus the assumption that the seeds of $D_{G}$ were sufficient to color all of $D_{G}$. We define a “double” of the original coloring move from Figure 4, which colors parallel strands one after the next, as shown in Figure 10. This is a special case of the original coloring move which is made out of repeated applications of the original move in a prescribed way. When coloring our link diagram $D_{L}$, we will restrict ourselves to only using this double coloring move, so that the coloring will directly correspond to a coloring in the graph diagram $D_{G}$. Thus we can color all of $D_{L}$ by finding a sequence of coloring moves in $D_{G}$ and applying the corresponding doubled coloring moves in $D_{L}$. Note that once all the free ends of a tree strand in the original diagram are colored, that is sufficient for all strands in the double to be colored; see Figure 11.
Figure 10: The “double” coloring move: if the overstrands are colored (not necessarily the same color) and two of the understrands on the same side of the crossing are colored (not necessarily the same color), then the move extends to the other understrands.
Figure 11: All free ends being colored induces the coloring of every strand in the double. (Note that here we depict all free ends and strands in the double as receiving the same color for the sake of simplicity, but in general this will not be the case.)
Thus we have a sufficient set of $2k$ seeds for $D_{L}$, so we may apply the methods of [Bla+20a, Theorem 1.3] to construct a smooth embedding of $L$ in $\mathbb{R}^{3}$ with exactly $2k$ local maxima, and conclude that the Wirtinger number of $L$ is bounded below by the bridge index of $L$. As before, we embed $G$ in $\mathbb{R}^{3}$ such that it projects to $D$ on the $xy$-plane, and adjust the height ($z$-coordinate) of pieces of the graph. We embed the seeds at the highest levels, and then embed the remaining strands in the order indicated by the coloring procedure. The result of [Bla+20a] ensures that we do not add new local maxima during the embedding process in addition to the seeds.
However, we also need to embed $L$ in such a way that respects the double structure, so we can ultimately translate back to our original graph. Thus as we embed we require that multiple strands which came from the same seed in $D_{G}$ must be embedded one after the other (in adjacent sheets), as depicted in Figure 12. Our double coloring move ensures that this will happen as the coloring progresses, so we need only add the requirement that the initial seeds are colored one after the other in this way.
Figure 12: We embed strands which came from the same seed in $D_{G}$ in adjacent sheets.
From this procedure we get an embedding of each strand of $D_{L}$ in a plane (parallel to the $xy$-plane) in $\mathbb{R}^{3}$. Now we connect up the strands in $\mathbb{R}^{3}$ to form a ribbon link whose core is our spatial graph. We then have an embedding of our spatial graph with bridge index $k$, since when we collapse the ribbon link to its core, any induced seeds in $D_{L}$ which came from the same seed in $D_{G}$ will collapse back to one seed. As we started with a Wirtinger number $k$ diagram, we can conclude that the Wirtinger number of $G$ is bounded below by the bridge index of $G$. ∎
Just as in [Bla+20a], our main result gives a way to pass between two classical ways to define the bridge index for links: one in terms of the number of overpasses and the other in terms of the number of local maxima. (Our definition for bridge index is in terms of intersections with the bridge sphere, but this is a rephrasing of the maxima definition necessitated by the setting of spatial graphs; see the comments before Definition 2.1.) We elaborate how to go between these formulations for spatial graphs now.
An overpass is a strand in a spatial graph diagram that contains at least one overcrossing (and by definition, contains no undercrossings). We define the overpass bridge index of a spatial graph diagram with overpasses $\{O_{1},\ldots,O_{n}\}$ to be $\frac{1}{2}\sum w_{i}$, where $w_{i}$ is the weight corresponding to the strand $O_{i}$. The overpass bridge index of a spatial graph $G$ is the minimum value of this sum over all diagrams representing $G$.
Corollary 3.3.
The overpass bridge index equals the bridge index.
Proof.
For a given spatial graph, take a diagram realizing the overpass bridge index and declare that the overpasses are seeds. The coloring can surely be extended to the entire diagram since all of the overcrossings are colored.
For the other inequality, take a bridge splitting realizing the bridge index. By definition, each trivial tangle is a collection of unknotted trees in a $3$-ball. This means that there exists a disk containing each unknotted tree that guides an isotopy of each tree down to the bridge splitting sphere so that at the end of the process, the trees in one trivial tangle $\tau_{1}$ lie completely in the bridge sphere without intersecting one another. Each component then corresponds to an overpass since performing the same process on the other trivial tangle $\tau_{2}$ gives another collection of trees on the bridge sphere, and we can declare that any intersection point in the diagram has an arc from $\tau_{1}$ crossing over an arc from $\tau_{2}.$
@@ 4 Lower bounds
The following techniques are computable quantities that bound the bridge index of graphs from below. In the setting of links, we know by experimentation that depending on the link types, one method may be more effective than the others [NPV24]. That is, in some instances the computer code finds enough quandle colorings, but not a rank bound, and vice versa.
In general, both techniques require an oriented spatial graph as input, meaning that we designate each edge as having a direction coming out of one vertex and into the other, denoted with an arrow. However we can disregard orientation when using involutory quandles (with techniques from Section 4.1) or groups where each generator in the label is its own inverse (with techniques from Section 4.2).
@@ 4.1 Quandle colorings
In [FM07, Nie10] the authors generalized knot quandles to spatial graphs.
@@
A consistent labeling [^2] Various authors also use the term quandle coloring, $X$-coloring, or simply coloring for this of a diagram by a quandle is an assignment of a quandle element to each arc, such that arcs connected to the same vertex receive the same label, and the relation in Figure 13 is satisfied at each crossing. It can be shown that for knots and links, the number of consistent labelings is unchanged by the Reidemeister moves. For general spatial graphs, however, this is no longer true.
@@
Figure 14: For a quandle coloring to work, the strands near a crossing need to all receive the same color or all distinct colors. On the left, the strands are disjoint and the arcs connected to the same vertex receive one color. If the other strand receives a different color, after we perform a move that does not change the graph type, we are forced to have a crossing that uses only two colors.
@@
This motivates the following definitions.
@@
Definition 4.3.
Give a finite $n$-quandle $X$ and a spatial graph $G$, the coloring number $Col_{X}(G)$ is the number of consistent labelings of $G$ by $X$.
Recall that when we compute the bridge index, which is shown to be equal to the Wirtinger number by Theorem 1.1, we assign each seed a weight. The traditional quandle counting invariant will give a bound on the number of seeds unweighted. When we apply it for practical purposes in Section 5, we will add in the weight consideration.
Definition 4.4.
Given a spatial graph $G$, define the unweighted bridge index $\widehat{\beta}(G)$ to be the minimum number of seeds over all diagrams of $G$ (with no weights assigned).
@@
Let $X$ be a finite $n$-quandle of order $|X|$. Then $Col_{X}(G)\leq|X|^{\widehat{\beta}(G)}$. In other words,
$\log_{|X|}Col_{X}(G)\leq\widehat{\beta}(G).$
@@
By Theorem 1.1, there exists a diagram $D$ realizing the $\widehat{\beta}(G)$ seeds. There are $|X|$ choices of quandle elements from $X$ to assign to each seed. The coloring move in Figure 4 can be regarded as a less restricted version of the quandle rule in Figure 13. More precisely, note that the rule in Figure 13 says that the outgoing quandle labeling is a result of using the quandle operation on the incoming arc and the overstrand, which is analogous to saying that the outgoing arc receives a coloring once the incoming arc and the overstrand receive colors in Figure 4. Thus, there are $|X|^{\widehat{\beta}(G)}$ possibilities of quandle labels for the seeds that will generate the quandle labeling for the entire diagram. The reason why $Col_{X}(G)\leq|X|^{\widehat{\beta}(G)}$ is possibly not an equality is because not every labeling generated from the seeds is consistent. ∎
We now provide a lower bound for the bridge index of vertex sums of spatial graphs (recall Section 2.1) in terms of quandles, which will be useful to us in Section 5. The argument is based on [CSV16]. The key is the existence of quandles called homogeneous quandles, which have the property that for any quandle elements $x,y,$ there is an automorphism $h$ such that $h(x)=y.$ Note that the map $f_{y}$ sending $x$ to $x\triangleright y$ in the second axiom of Definition 4.1 is an example of an automorphism by the way the axiom is stated. The reader can check from the table before Definition 4.2 that the Alexander quandle of order four is homogeneous.
@@
Let $X$ be a homogeneous $n$-quandle, and $G\#_{n}G^{\prime}$ be the $n$-valent vertex sum of spatial graphs $G$ and $G^{\prime}$. Then
@@
$\log_{|X|}Col_{X}(G)+\log_{|X|}Col_{X}(G^{\prime})-1\leq\widehat{\beta}(G\#_{n}G^{\prime}).$
@@
The number of colorings where a fixed color $x$ is assigned to the arcs adjacent to a vertex $v$ is precisely $Col_{X}(G)/|X|$. To see this, observe that the number of colorings $Col_{X}(G)=\sum_{y}Col_{X}(G,y)$, where $Col_{X}(G,y)$ denotes the number of colorings such that the arcs near $v$ get the label $y$. We can get another coloring in which the arcs adjacent to $v$ are colored $z\neq y$ by applying an automorphism $h$ such that $h(y)=z,$ which exists by the homogeneity assumption. Thus, $Col_{X}(G)=|X|Col_{X}(G,y)$.
@@
The condition that the colors of the arcs adjacent to a vertex have to be the same can be relaxed. In fact, this may be needed to bound our definition of the bridge index in which we count the number of intersections of the spatial graph with bridge spheres (see Definition 2.1 and the following discussion). This is pointed out in [IY97] with a brief argument and a reference to the proof in a different paper. [^3] To the best of the authors’ knowledge, this other paper never appeared in the literature. We provide the proof again here for accessibility and completeness, as some examples in Section 5 will need this type of coloring.
Definition 4.7.
A generalized $3$-coloring of a $4$-regular spatial graph diagram is an assignment of three elements $\{0,1,2\}$ to the arcs such that the following two properties hold.
  1. 1.
    If $x,y,$ and $z$ are the incoming arc, outgoing arc, and the overstrand at a crossing respectively, then $2z-y=x\pmod{3}$.
  2. 2.
    Let $a_{1},a_{2},a_{3},a_{4}$ be labels around a vertex arranged in a counterclockwise fashion. Then $a_{1}-a_{2}+a_{3}-a_{4}=0\pmod{3}$.
Proposition 4.8.
The number of generalized $3$-colorings of a $4$-regular spatial graph diagram is a spatial graph invariant.
Proof.
The condition around each crossing is satisfied because tricolorability is known to be a knot invariant. At a vertex, we have to check that there is a bijection before and after Reidemeister moves. In Figure 15, all the Reidemeister moves involving vertices are depicted. The labels after a chain of quandle rules for the top left picture will give y=2ax,z=2by,w=2cz,y=2a-x,z=2b-y,w=2c-z, and t=2dw.t=2d-w. In other words, t=2d(2c(2b(2ax)))=2(ab+cd)+x=xt=2d-(2c-(2b-(2a-x)))=-2(a-b+c-d)+x=x since $a-b+c-d=0\pmod{3}$. For the other version of the same move, we have that e=2xa,f=2xb,g=2xc,e=2x-a,f=2x-b,g=2x-c, and h=2xdh=2x-d. Since we have $e-f+g-h=(2x-a)-(2x-b)+(2x-c)-(2x-d)=0\pmod{3}$, the coloring is invariant.
For the twisting move at a vertex, we check that the sum ji+lr=0,j-i+l-r=0, but r=2lkr=2l-k so that ji+l2l+k=jil+k=0.j-i+l-2l+k=j-i-l+k=0.
[…]
Figure 15: The number of generalized tricolorings is preserved by Reidemeister moves involving vertices.
Remark 4.9.
McAtee, Silver, and Williams consider colorings of spatial graphs using topological groups in [MSW01]. Since any topological group is a quandle [Rub07], it would be interesting to combine the quandle techniques discussed in this section with the techniques of [MSW01].